OpenStudy (anonymous):

Let f(x)={4x−1, if x≤3...... .............−5x+b, if x>3 If f(x) is a function which is continuous everywhere, then we must have b=

myininaya (myininaya):

do the left limit and right limit for x=3 and set them equal and solve for b

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

how would i go about solving this one...same way? For what value of the constant c is the function f continuous on (−∞,∞) where f(x)=...... cx+8 if x∈(−∞,2] ....... cx^2−8 if x∈(2,∞)......

myininaya (myininaya):

yes you look at the left limit and right limit for x=2

myininaya (myininaya):

and set them equal

OpenStudy (anonymous):

hey so for the first problem...i used 4(3)-1 = 11 and -5(3.1)+b = -15.5 + b. So I set -15.5 + b = 11 and I did the algebra and ended up with -4.5. what did i do wrong?

myininaya (myininaya):

\[4(3)-1=-5(3)+b\]

OpenStudy (anonymous):

yeah shouldnt u get a -4.5 after you do the algebra

myininaya (myininaya):

i don't know i can perform the operations if you want me too

myininaya (myininaya):

4(3)=12 12-1=11 -5(3)=-15 so we have 11=-15+b By adding 15 on both sides we get 26=b

OpenStudy (anonymous):

oh snap i did it backward haha thanks

OpenStudy (anonymous):

can i use 1.5 and 3 for the second question

myininaya (myininaya):

i don't understand

OpenStudy (anonymous):

how would i go about solving this one...same way? For what value of the constant c is the function f continuous on (−∞,∞) where f(x)=...... cx+8 if x∈(−∞,2] ....... cx^2−8 if x∈(2,∞)...... once again i have to approach from both sides...if im approaching from left can I use x=1.5 and right x=3?

myininaya (myininaya):

what no just before we plugged in 3 for both sides here we do 2

myininaya (myininaya):

\[c(2)+8=c(2)^2-8\]

myininaya (myininaya):

now solve for c